OpenStudy (anonymous):

Which is an equation of the line passing through points at (1, 6) and (-2, 1)?

OpenStudy (anonymous):

answer choices y = 2x + 5 y = 5/3x + 13/3 y = x + 5 y = -1/2x +13/2

OpenStudy (mathteacher1729):

Do you know how to find the slope of a line given two points?

OpenStudy (anonymous):

naww i just need help finding the answer is it y= mx+b?

OpenStudy (mathteacher1729):

y = mx + b is the equation of a straight line with slop = m and y-intercept = b.

OpenStudy (mathteacher1729):

The slope is found by taking the "rise over run" or "difference is y over difference in x".

OpenStudy (anonymous):

I am not good at math so i have no idea what you just said. lol :/ could you please just tell me the answer?

OpenStudy (mathteacher1729):

Sometimes a picture is worth a thousand words. 1) Plot the two points. 2) Find out how far "up" you go 3) Find out how far "right" you go. In this case up = rise = 5 and right = run = 3 so the slop is just 5 / 3.

OpenStudy (anonymous):

okay gotcha. so would the answer be y = 5/3x + 13/3 ?

OpenStudy (mathteacher1729):

*slope not "slop" lol.

OpenStudy (mathteacher1729):

Well, it's the only option with a slope of 5/3. :) Just check to see if 13/3 is really 4.33333 and you're done.

OpenStudy (anonymous):

yes it is! lol thank you! could you help me with some other problems?

OpenStudy (mathteacher1729):

I'm going to sleep for the evening. Got a long day tomorrow. :) I'm glad this was helpful though. The program I used is called Geogebra. http://www.geogebra.org/cms/en/installers I HIGHLY recommend it, because it can solve problems like this in about 2 seconds. :)

OpenStudy (anonymous):

ok thank you